# LoRA: Low-Rank Adaptation of Large Language Models

### v1

@@ 4 Related Works
Transformer Language Models.  Transformer [36] is a sequence-to-sequence architecture that makes heavy use of self-attention. [31] applied it to autoregressive language modeling by using a stack of Transformer decoders. Since then, Transformer-based language models have dominated NLP, achieving the state-of-the-art in many tasks. A new paradigm emerged with BERT [9] and GPT-2 [32] – both are large Transformer language models trained on a large amount of text – where fine-tuning on task-specific data after pre-training on general domain data provides a significant performance gain compared to training on task-specific data directly. Training larger Transformers generally results in better performance and remains an active research direction. GPT-3 [6] is the largest single Transformer language model trained to-date with 175B parameters.
@@
Prompt Engineering and Fine-Tuning.  While GPT-3 175B can adapt its behavior with just a few additional training examples, the result depends heavily on the input prompt [6]. This necessitates an empirical art of composing and formatting the prompt to maximize a model’s performance on a desired task, which is known as prompt engineering or prompt hacking. Fine-tuning retrains a model pre-trained on general domains to a specific task [9, 31]. Variants of it include learning just a subset of the parameters [9, 8], yet practitioners often retrain all of them to maximize the downstream performance. However, the enormity of GPT-3 175B makes it challenging to perform fine-tuning in the usual way due to the large checkpoint it produces and the high hardware barrier to entry since it has the same memory footprint as pre-training.
@@
Parameter-Efficient Adaptation.  [15, 34] propose inserting adapter layers between existing layers in a neural network. Our method uses a bottleneck structure similar to [15] to impose a low-rank constraint on the weight updates. The key functional difference is that our learned weights can be merged with the main weights during inference, thus not introducing any latency, which is not the case for the adapter layers. More recently, [21, 19, 14, 25] proposed optimizing the input word embeddings in lieu of fine-tuning, akin to a continuous and differentiable generalization of prompt engineering. We include comparisons with [21] in our experiment section. However, this line of works can only scale up by using more special tokens in the prompt, which take up available sequence length for task tokens when positional embeddings are learned.
@@
Low-Rank Structures in Deep Learning.  Low-rank structure is very common in machine learning. A lot of machine learning problems have certain intrinsic low-rank structure [23, 7, 24, 13]. Moreover, it is known that for many deep learning tasks, especially those with a heavily over-parametrized neural network, the learned neural network will enjoy low-rank properties after training [29]. Some prior works even explicitly impose the low-rank constraint when training the original neural network, such as [35, 30, 38, 17, 39], however, to the best of our knowledge, none of these works considers low-rank update for adaptation to downstream tasks. In theory literature, it is known that neural networks outperform other classical learning methods, including the corresponding (finite-width) neural tangent kernels [5, 22] when the underlying concept class has certain low-rank structure [11, 2, 3]. Another theoretical result in [4] suggests that low-rank adaptations can be useful for adversarial training. In sum, we believe that our proposed low-rank adaptation update is well-motivated by the literature.
5 Empirical Experiments
We benchmark the downstream performance of LoRA on both GPT-2 and GPT-3. Specifically, we focus on WikiSQL [40] (natural language to SQL queries), MultiNLI [37] (natural language inference)  [^3] Recent works treat MultiNLI-matched and MultiNLI-mismatched as separate tasks. In this work, we only report accuracy on MultiNLI-matched validation set., and SAMSum [12] (conversation summarization) for GPT-3.For a direct comparison with  [21] on GPT-2, we follow their setup and use Table-to-Text E2E NLG Challenge  [28], a language generation task. See Appendix B for more details on these datasets. We also compare with the other tuning methods such as [21, 25]. We use NVIDIA Tesla V100 GPUs for all of our experiments. For GPT-3, we use 96 GPUs for fine-tuning and 24 GPUs for LoRA. We use a single GPU for GPT-2 experiments.
5.1 Baselines
We compare with the following baselines. Note that we implemented both prefix-embedding and prefix-layer tuning in GPT-3 on our own according to [21], while the non-LoRA baselines all come from the implementation of [21].
Fine-Tuning is the de facto default adaptation approach, where the model is initialized to the pre-trained parameters. During fine-tuning, all the model parameters undergo gradient updates according to Eqn. 1. A simple variant of it is to update only some layers while freezing others. We include one such baseline reported in prior work [21], which tunes just the last two layers (FT-Top2).
Bias only is a baseline where we only train the bias vectors when adapting to downstream tasks.
Prefix-embedding tuning injects special tokens alongside the input tokens. These special tokens have trainable word embeddings and are generally not in the model’s vocabulary. The placement of such tokens can have an impact on performance. We focus on “prefixing”, which prepends these tokens to the prompt, and “infixing”, which appends to the prompt; both are discussed in [21]. We use $l_{p}$ (resp. $l_{i}$) denote the number of prefix (resp. infix) tokens. The number of trainable parameters is $|\Theta|=d_{model}\times(l_{p}+l_{i})$.
Prefix-layer tuning is an extension to prefix-embedding tuning. Instead of just learning the word embeddings (or equivalently, the activations after the embedding layer) for some special tokens, we learn the activations after every layer. The activations computed from previous layers are simply replaced by our trainable ones. The resulting number of trainable parameters is $|\Theta|=L\times d_{model}\times(l_{p}+l_{i})$, where $L$ is the number of layers (96 for GPT-3).
Adapter tuning [15] inserts adapter layers between the self-attention module (and the MLP module) and the subsequent residual connection. There are two weight matrices in an adapter layer with a nonlinearity in between. The dimension of the hidden layer in an adapter layer determines the total number of trainable parameters. For a hidden size of n, we have $|\Theta|=4\times L\times d_{model}\times n$.
LoRA adds trainable pairs of rank decomposition matrices to existing weight matrices. As mentioned in Sec. 3.1 and supported by our analysis in Sec. 6.1, we only apply LoRA to $W_{q}$ and $W_{v}$ in our experiments. The number of trainable parameters is determined by the rank $r$ and the shape of the original weights: $|\Theta|=2\times L\times d_{model}\times r$.
Method # of Trainable WikiSQL MNLI-m SAMSum
Parameters Accuracy (%) Accuracy (%) R1/R2/RL
GPT-3 175B (Fine-Tune) 175,255.8M 73.0 89.5 52.0/28.0/44.5
GPT-3 175B (Bias Only) 14.2M 71.3 91.0 51.3/27.4/43.5
GPT-3 175B (PrefixEmbed) 3.2M 63.1 88.6 48.3/24.2/40.5
GPT-3 175B (PrefixLayer) 20.2M 70.1 89.5 50.8/27.3/43.5
GPT-3 175B (LoRA) 4.7M 73.4 91.3 52.1/28.3/44.3
GPT-3 175B (LoRA) 37.7M 73.8 91.7 53.2/29.2/45.0
Table 1: Logical form validation accuracy on WikiSQL, validation accuracy on MultiNLI-matched and Rouge-1/2/L on SAMSum achieved by different GPT-3 adaptation methods. LoRA performs better than prior approaches, including conventional fine-tuning. The result on WikiSQL has a fluctuation of $\pm 0.3\%$ and MNLI-m $\pm 0.1\%$.
5.2 Performance on GPT-3
Hyperparameters
For all of our GPT-3 experiments, we train using AdamW [26] for 2 epochs with a batch size of 100k tokens, a sequence length of 768, and a weight decay factor of 0.1. We tune the learning rate for all method-dataset combinations. See Sec. C.1 for more details. For prefix-embedding tuning, we find the optimal $l_{p}$ and $l_{i}$ to be 256 and 8, respectively, totalling $3.2M$ trainable parameters. We use $l_{p}=8$ and $l_{i}=8$ in PrefixLayer with $20.2M$ trainable parameters to obtain the overall best performance. We present two configurations for LoRA: one with $r_{q}=r_{v}=4$ (18.8M) and one with $r_{q}=r_{v}=8$ (37.7M).
As shown in Table 1, on all three datasets, LoRA outperforms the fine-tuning baseline. It might appear that prefix-embedding tuning could benefit from having more trainable parameters since it uses much fewer parameters than other methods. Nonetheless, this is not the case as shown in Fig. 2. We observe a significant performance drop when we use more than 256 special tokens for prefix-embedding tuning or more than 32 special tokens for prefix-layer tuning. This corroborates the similar observation reported in [21]. While a thorough investigation into this phenomenon is out-of-scope for this work, we suspect that having more special tokens causes the input distribution to shift further away from the pre-training data distribution, degrading the model performance. Separately, we investigate the performance of adaptation approaches in the low-data regime in Sec. E.3.
Figure 2: GPT-3 175B validation accuracy vs. number of trainable parameters of several adaptation methods on WikiSQL and MNLI-matched. LoRA enjoys better scalability and task performance. See Sec. E.1 for more details on the plotted data points.
5.3 Performance on GPT-2
Having established LoRA as a competitive method on GPT-3, we hope to answer if LoRA can also work on less overly parametrized models, such as GPT-2 medium or large [32]. We keep our setup as close as possible to [21] for a direct comparison. Due to space constraint, we only present our result on E2E NLG Challenge (Table 2) in this section – see Sec. E.2 for results on more datasets.
Hyperparameters
We train all of our GPT-2 models using AdamW [26] and a linear learning rate schedule for 5 epochs. We use the batch size, learning rate, and beam search beam size described in [21]. Accordingly, we also tune these hyperparameters for LoRA, and we include a list of the used hyperparameters in Table 8.
Method # of Trainable E2E
Parameters BLEU NIST MET ROUGE-L CIDEr
GPT-2 M (Fine-Tune) 354.92M 68.2 8.62 46.2 71.0 2.47
GPT-2 M (Adapter) 11.48M 68.9 8.71 46.1 71.3 2.47
GPT-2 M (FT-Top2) 25.19M 68.1 8.59 46.0 70.8 2.41
GPT-2 M (Prefix) 0.35M 69.7 8.81 46.1 71.4 2.49
GPT-2 M (LoRA) 0.35M 70.4 8.85 46.8 71.8 2.53
GPT-2 L (Fine-Tune) 774.03M 68.5 8.78 46.0 69.9 2.45
GPT-2 L (Prefix) 0.77M 70.3 8.85 46.2 71.7 2.47
GPT-2 L (LoRA) 0.77M 70.4 8.89 46.8 72.0 2.47
Table 2: GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge. For all metrics, higher is better. LoRA outperforms several baselines with comparable or fewer trainable parameters.
@@ 6 Understanding the Low-Rank Updates
Given the empirical advantage of LoRA, we hope to further explain the properties of the low-rank adaptation learned from downstream tasks. Note that the low-rank structure not only lowers the hardware barrier to entry which allows us to run multiple experiments in parallel, but also gives better interpretability of how the update weights are correlated with the pre-trained weights.
@@ 6.1 Which Weight Matrices in Transformer Should We Apply LoRA to?
Given a limited parameter budget, which types of weights should we adapt with LoRA to obtain the best performance on downstream tasks? As mentioned in Sec. 3.1, we only consider weight matrices in the self-attention module. We set a parameter budget of 18M (roughly 35MB) on GPT-3, which corresponds to $r=8$ if we adapt one type of weights or $r=4$ if we adapt two types, for all 96 layers. The result is presented in Table 3.
@@
Weight Type $W_{q}$ $W_{k}$ $W_{v}$ $W_{o}$ $W_{q},W_{k}$ $W_{q},W_{v}$
Rank $r$ 8 8 8 8 4 4
WikiSQL ($\pm 0.3$%) 70.4 70.0 73.0 73.2 71.4 73.7
MultiNLI ($\pm 0.1$%) 91.0 90.8 91.0 91.3 91.3 91.3
@@ 6.2 What is the Optimal Rank $r$ for LoRA?
We turn our attention to the effect of rank $r$ on model performance. We adapt both $W_{q}$ and $W_{v}$ since they performed the best in the previous experiment and as well as $W_{q}$ alone for comparison.
@@
WikiSQL($\pm 0.3$%) $W_{q},W_{v}$ 73.4 73.3 73.7 73.8 73.5
$W_{q}$ 68.8 69.6 70.5 70.4 70.0
MultiNLI ($\pm 0.1$%) $W_{q},W_{v}$ 91.3 91.4 91.3 91.7 91.4
Table 4: Validation accuracy on WikiSQL and MultiNLI with different rank $r$. To our surprise, a rank as small as one suffices for adapting both $W_{q}$ and $W_{v}$ on these datasets while training $W_{q}$ alone needs a larger $r$. We replicate this on GPT-2 in Sec. G.2.
Table 4 shows that, surprisingly, LoRA already performs competitively with a very small $r$ (more so for $W_{q+v}$ than $W_{q}$). This suggests the update matrix $\Delta W$ could have a very small “intrinsic rank". [^4] However, we do not expect a small $r$ to work for every task or dataset. Consider the following thought experiment: if the downstream task were in a different language than the one used for pre-training, retraining the entire model (similar to LoRA with $r=d_{model}$) could certainly outperform LoRA with a small $r$.
To further support this finding, we check the overlap of the subspaces learned by different choices of $r$ and by different random seeds. We argue that increasing $r$ does not cover more meaningful subspaces, which suggests that a low-rank adaptation matrix is sufficient.
Subspace similarity between different $r$.  Given $A_{r=8}$ and $A_{r=64}$ which are the learned adaptation matrices with rank $r=8$ and $64$ using the same pre-trained model, we perform singular value decomposition and obtain the right-singular unitary matrices $U_{A_{r=8}}$ and $U_{A_{r=64}}$. [^5] Note that a similar analysis can be carried out with $B$ and the left-singular unitary matrices – we stick with $A$ for our experiments. We hope to answer: how much of the subspace spanned by the top $i$ singular vectors in $U_{A_{r=8}}$ (for $1\leq i\leq 8$) is contained in the subspace spanned by top $j$ singular vectors of $U_{A_{r=64}}$ (for $1\leq j\leq 64$)? We measure this quantity with a normalized subspace similarity based on the Grassmann distance (See Appendix F for a more formal discussion)
@@
$\phi(\cdot)$ has a range of $[0,1]$, where $1$ represents a complete overlap of subspaces and 0 a complete separation. See Fig. 3 for how $\phi$ changes as we vary $i$ and $j$. We only look at the 48th layer (out of 96) due to space constraint, but the conclusion holds for other layers as well, as shown in Sec. G.1.
@@
We make an important observation from Fig. 3.
@@
Since both $A_{r=8}$ and $A_{r=64}$ are learned using the same pre-trained model, Fig. 3 indicates that the top singular-vector directions of $A_{r=8}$ and $A_{r=64}$ are the most useful, while other directions potentially contain mostly random noises accumulated during training. Hence, the adaptation matrix can indeed have a very low rank.
Figure 4: Left and Middle: Normalized subspace similarity between the column vectors of $A_{r=64}$ from two random seeds, for both $\Delta W_{q}$ and $\Delta W_{v}$ in the 48-th layer. Right: the same heat-map between the column vectors of two random Gaussian matrices. See Sec. G.1 for other layers.
Subspace similarity between different random seeds.  We further confirm this by plotting the normalized subspace similarity between two randomly seeded runs with $r=64$, shown in Fig. 4. $\Delta W_{q}$ appears to have a higher “intrinsic rank” than $\Delta W_{v}$, since more common singular value directions are learned by both runs for $\Delta W_{q}$, which is in line with our empirical observation in Table 4. As a comparison, we also plot two random Gaussian matrices, which do not share any common singular value directions with each other.
@@ 6.3 How Does the Adaptation Matrix $\Delta W$ Compare to $W$?
We draw several conclusions from Table 5. First, $\Delta W$ has a stronger correlation with $W$ compared to a random matrix, indicating that $\Delta W$ amplifies some features that are already in $W$. Second, instead of repeating the top singular directions of $W$, $\Delta W$ only amplifies directions that are not emphasized in $W$. Third, the amplification factor is rather huge: $21.5\approx 6.91/0.32$ for $r=4$. See Sec. G.4 for why $r=64$ has a smaller amplification factor. We also provide a visualization in Sec. G.3 for how the correlation changes as we include more top singular directions from $W_{q}$. This suggests that the low-rank adaptation matrix potentially amplifies the important features for specific downstream tasks that were learned but not emphasized in the general pre-training model.

### v2

@@ 6 Related Works
Transformer Language Models.  Transformer (Vaswani et al., 2017) is a sequence-to-sequence architecture that makes heavy use of self-attention. Radford et al. (a) applied it to autoregressive language modeling by using a stack of Transformer decoders. Since then, Transformer-based language models have dominated NLP, achieving the state-of-the-art in many tasks. A new paradigm emerged with BERT (Devlin et al., 2019b) and GPT-2 (Radford et al., b) – both are large Transformer language models trained on a large amount of text – where fine-tuning on task-specific data after pre-training on general domain data provides a significant performance gain compared to training on task-specific data directly. Training larger Transformers generally results in better performance and remains an active research direction. GPT-3 (Brown et al., 2020) is the largest single Transformer language model trained to-date with 175B parameters.
@@
Prompt Engineering and Fine-Tuning.  While GPT-3 175B can adapt its behavior with just a few additional training examples, the result depends heavily on the input prompt (Brown et al., 2020). This necessitates an empirical art of composing and formatting the prompt to maximize a model’s performance on a desired task, which is known as prompt engineering or prompt hacking. Fine-tuning retrains a model pre-trained on general domains to a specific task Devlin et al. (2019b); Radford et al. (a). Variants of it include learning just a subset of the parameters Devlin et al. (2019b); Collobert & Weston (2008), yet practitioners often retrain all of them to maximize the downstream performance. However, the enormity of GPT-3 175B makes it challenging to perform fine-tuning in the usual way due to the large checkpoint it produces and the high hardware barrier to entry since it has the same memory footprint as pre-training.
@@
Parameter-Efficient Adaptation.  Many have proposed inserting adapter layers between existing layers in a neural network (Houlsby et al., 2019; Rebuffi et al., 2017; Lin et al., 2020). Our method uses a similar bottleneck structure to impose a low-rank constraint on the weight updates. The key functional difference is that our learned weights can be merged with the main weights during inference, thus not introducing any latency, which is not the case for the adapter layers (Section 3). A comtenporary extension of adapter is compacter (Mahabadi et al., 2021), which essentially parametrizes the adapter layers using Kronecker products with some predetermined weight sharing scheme. Similarly, combining LoRA with other tensor product-based methods could potentially improve its parameter efficiency, which we leave to future work. More recently, many proposed optimizing the input word embeddings in lieu of fine-tuning, akin to a continuous and differentiable generalization of prompt engineering (Li & Liang, 2021; Lester et al., 2021; Hambardzumyan et al., 2020; Liu et al., 2021). We include comparisons with Li & Liang (2021) in our experiment section. However, this line of works can only scale up by using more special tokens in the prompt, which take up available sequence length for task tokens when positional embeddings are learned.
@@
Low-Rank Structures in Deep Learning.  Low-rank structure is very common in machine learning. A lot of machine learning problems have certain intrinsic low-rank structure (Li et al., 2016; Cai et al., 2010; Li et al., 2018b; Grasedyck et al., 2013). Moreover, it is known that for many deep learning tasks, especially those with a heavily over-parametrized neural network, the learned neural network will enjoy low-rank properties after training (Oymak et al., 2019). Some prior works even explicitly impose the low-rank constraint when training the original neural network (Sainath et al., 2013; Povey et al., 2018; Zhang et al., 2014; Jaderberg et al., 2014; Zhao et al., 2016; Khodak et al., 2021; Denil et al., 2014); however, to the best of our knowledge, none of these works considers low-rank update to a frozen model for adaptation to downstream tasks. In theory literature, it is known that neural networks outperform other classical learning methods, including the corresponding (finite-width) neural tangent kernels (Allen-Zhu et al., 2019; Li & Liang, 2018) when the underlying concept class has certain low-rank structure (Ghorbani et al., 2020; Allen-Zhu & Li, 2019; 2020a). Another theoretical result in Allen-Zhu & Li (2020b) suggests that low-rank adaptations can be useful for adversarial training. In sum, we believe that our proposed low-rank adaptation update is well-motivated by the literature.
@@ 7 Understanding the Low-Rank Updates
[…]
@@ 7.1 Which Weight Matrices in Transformer Should We Apply LoRA to?
Given a limited parameter budget, which types of weights should we adapt with LoRA to obtain the best performance on downstream tasks? As mentioned in Section 4.2, we only consider weight matrices in the self-attention module. We set a parameter budget of 18M (roughly 35MB if stored in FP16) on GPT-3 175B, which corresponds to $r=8$ if we adapt one type of attention weights or $r=4$ if we adapt two types, for all 96 layers. The result is presented in Table 5.
@@
Weight Type $W_{q}$ $W_{k}$ $W_{v}$ $W_{o}$ $W_{q},W_{k}$ $W_{q},W_{v}$ $W_{q},W_{k},W_{v},W_{o}$
Rank $r$ 8 8 8 8 4 4 2
WikiSQL ($\pm 0.5$%) 70.4 70.0 73.0 73.2 71.4 73.7 73.7
MultiNLI ($\pm 0.1$%) 91.0 90.8 91.0 91.3 91.3 91.3 91.7
@@ 7.2 What is the Optimal Rank $r$ for LoRA?
We turn our attention to the effect of rank $r$ on model performance. We adapt $\{W_{q},W_{v}\}$, $\{W_{q},W_{k},W_{v},W_{c}\}$, and just $W_{q}$ for a comparison.
@@
WikiSQL($\pm 0.5$%) $W_{q}$ 68.8 69.6 70.5 70.4 70.0
$W_{q},W_{v}$ 73.4 73.3 73.7 73.8 73.5
$W_{q},W_{k},W_{v},W_{o}$ 74.1 73.7 74.0 74.0 73.9
MultiNLI ($\pm 0.1$%) $W_{q}$ 90.7 90.9 91.1 90.7 90.7
$W_{q},W_{v}$ 91.3 91.4 91.3 91.6 91.4
$W_{q},W_{k},W_{v},W_{o}$ 91.2 91.7 91.7 91.5 91.4
Table 6: Validation accuracy on WikiSQL and MultiNLI with different rank $r$. To our surprise, a rank as small as one suffices for adapting both $W_{q}$ and $W_{v}$ on these datasets while training $W_{q}$ alone needs a larger $r$. We conduct a similar experiment on GPT-2 in Section H.2.
Table 6 shows that, surprisingly, LoRA already performs competitively with a very small $r$ (more so for $\{W_{q},W_{v}\}$ than just $W_{q}$). This suggests the update matrix $\Delta W$ could have a very small “intrinsic rank”. [^6] However, we do not expect a small $r$ to work for every task or dataset. Consider the following thought experiment: if the downstream task were in a different language than the one used for pre-training, retraining the entire model (similar to LoRA with $r=d_{model}$) could certainly outperform LoRA with a small $r$. To further support this finding, we check the overlap of the subspaces learned by different choices of $r$ and by different random seeds. We argue that increasing $r$ does not cover a more meaningful subspace, which suggests that a low-rank adaptation matrix is sufficient.
Subspace similarity between different $r$.  Given $A_{r=8}$ and $A_{r=64}$ which are the learned adaptation matrices with rank $r=8$ and $64$ using the same pre-trained model, we perform singular value decomposition and obtain the right-singular unitary matrices $U_{A_{r=8}}$ and $U_{A_{r=64}}$. [^7] Note that a similar analysis can be carried out with $B$ and the left-singular unitary matrices – we stick with $A$ for our experiments. We hope to answer: how much of the subspace spanned by the top $i$ singular vectors in $U_{A_{r=8}}$ (for $1\leq i\leq 8$) is contained in the subspace spanned by top $j$ singular vectors of $U_{A_{r=64}}$ (for $1\leq j\leq 64$)? We measure this quantity with a normalized subspace similarity based on the Grassmann distance (See Appendix G for a more formal discussion)
@@
$\phi(\cdot)$ has a range of $[0,1]$, where $1$ represents a complete overlap of subspaces and 0 a complete separation. See Figure 3 for how $\phi$ changes as we vary $i$ and $j$. We only look at the 48th layer (out of 96) due to space constraint, but the conclusion holds for other layers as well, as shown in Section H.1.
@@
We make an important observation from Figure 3.
@@
Since both $A_{r=8}$ and $A_{r=64}$ are learned using the same pre-trained model, Figure 3 indicates that the top singular-vector directions of $A_{r=8}$ and $A_{r=64}$ are the most useful, while other directions potentially contain mostly random noises accumulated during training. Hence, the adaptation matrix can indeed have a very low rank.
Figure 4: Left and Middle: Normalized subspace similarity between the column vectors of $A_{r=64}$ from two random seeds, for both $\Delta W_{q}$ and $\Delta W_{v}$ in the 48-th layer. Right: the same heat-map between the column vectors of two random Gaussian matrices. See Section H.1 for other layers.
Subspace similarity between different random seeds.  We further confirm this by plotting the normalized subspace similarity between two randomly seeded runs with $r=64$, shown in Figure 4. $\Delta W_{q}$ appears to have a higher “intrinsic rank” than $\Delta W_{v}$, since more common singular value directions are learned by both runs for $\Delta W_{q}$, which is in line with our empirical observation in Table 6. As a comparison, we also plot two random Gaussian matrices, which do not share any common singular value directions with each other.
@@ 7.3 How Does the Adaptation Matrix $\Delta W$ Compare to $W$?
We draw several conclusions from Table 7. First, $\Delta W$ has a stronger correlation with $W$ compared to a random matrix, indicating that $\Delta W$ amplifies some features that are already in $W$. Second, instead of repeating the top singular directions of $W$, $\Delta W$ only amplifies directions that are not emphasized in $W$. Third, the amplification factor is rather huge: $21.5\approx 6.91/0.32$ for $r=4$. See Section H.4 for why $r=64$ has a smaller amplification factor. We also provide a visualization in Section H.3 for how the correlation changes as we include more top singular directions from $W_{q}$. This suggests that the low-rank adaptation matrix potentially amplifies the important features for specific downstream tasks that were learned but not emphasized in the general pre-training model.